# GCRPNet: Graph-Enhanced Contextual and Regional Perception Network for Salient Object Detection in Optical Remote Sensing Images

### v3

@@ III-E Loss Function
IV EXPERIMENTS
IV-A Experimental Setup
IV-A1 Datasets
To evaluate the effectiveness of the method, we trained and tested it on two public benchmark datasets for ORSI-SOD: ORSSD[29] and EORSSD[30].
@@ IV-A1 Datasets
IV-A2 Implementation Details
In this study, we implemented the proposed method using the PyTorch framework and conducted all experiments on an RTX 4090Ti GPU. We set the learning rate to 1e-4, the batch size to 4, and employed the AdamW optimization strategy for training. To ensure consistency in the input data, each image was resized to 384×384 pixels before being fed into the model. Additionally, to effectively mitigate overfitting during training, we introduced a series of data augmentation techniques, including random scaling and cropping, random image transformations, and normalization. Finally, we trained the model for 100 epochs on the ORSSD and EORSSD datasets.
TABLE I: Quantitative comparison results in terms of $F_{\beta}^{max}$, $F_{\beta}^{mean}$, $F_{\beta}^{adp}$, $E_{\xi}^{max}$, $E_{\xi}^{mean}$, $E_{\xi}^{adp}$, $S_{\alpha}$ and MAEMAE score are presented on two benchmark datasets. In the tables, $\uparrow$ and $\downarrow$ indicate that higher and lower values are preferred, respectively. The top three results in each column are highlighted in red, blue, and green, respectively.
Methods Pubs EORSSD ORSSD
$F_{\beta}^{max}\uparrow$ $F_{\beta}^{mean}\uparrow$ $F_{\beta}^{adp}\uparrow$ $E_{\xi}^{max}\uparrow$ $E_{\xi}^{mean}\uparrow$ $E_{\xi}^{adp}\uparrow$ $S_{\alpha}\uparrow$ $MAE\downarrow$ $F_{\beta}^{max}\uparrow$ $F_{\beta}^{mean}\uparrow$ $F_{\beta}^{adp}\uparrow$ $E_{\xi}^{max}\uparrow$ $E_{\xi}^{mean}\uparrow$ $E_{\xi}^{adp}\uparrow$ $S_{\alpha}\uparrow$ $MAE\downarrow$
EMFINet TGRS’21 0.8721 0.8488 0.7987 0.9712 0.9603 0.9502 0.9292 0.0084 0.9002 0.8856 0.8617 0.9737 0.9671 0.9663 0.9366 0.0109
DAFNet TIP’21 0.8735 0.7982 0.6525 0.9816 0.9382 0.8216 0.9186 0.0053 0.8999 0.8442 0.7565 0.9821 0.9536 0.8928 0.9188 0.0106
ERPNet TCYB’22 0.8744 0.8272 0.7174 0.9665 0.9364 0.8995 0.9254 0.0082 0.9036 0.8799 0.8392 0.9738 0.9604 0.9531 0.9352 0.0114
SBA-Net ISPRS’22 0.8164 0.7984 0.7104 0.9382 0.9235 0.8704 0.8612 0.0142 0.8373 0.8246 0.8159 0.9394 0.9205 0.9286 0.8618 0.0264
HFANet TGRS’22 0.8950 0.8805 0.8651 0.9682 0.9608 0.9461 0.9298 0.0127 0.9059 0.8917 0.8697 0.9727 0.9662 0.9467 0.9355 0.0116
ESGNet TCSVT’23 0.8907 0.8754 0.8634 0.9781 0.9732 0.9462 0.9374 0.0051 0.9220 0.9113 0.9076 0.9858 0.9814 0.9661 0.9484 0.0065
ACCoNet TCYB’23 0.8822 0.8464 0.7787 0.9759 0.9659 0.9381 0.9303 0.0067 0.9149 0.8952 0.8830 0.9819 0.9775 0.9770 0.9428 0.0087
BSCGNet TGRS’23 0.8809 0.8666 0.8506 0.9693 0.9650 0.9644 0.9296 0.0079 0.9101 0.8977 0.8897 0.9784 0.9744 0.9773 0.9411 0.0093
SFANet TGRS’24 0.8834 0.8682 0.8495 0.9770 0.9725 0.9682 0.9351 0.0058 0.9192 0.9063 0.8984 0.9830 0.9789 0.9800 0.9453 0.0077
RAGRNet TGRS’24 0.8853 0.8771 0.8646 0.9785 0.9736 0.9761 0.9364 0.0057 0.9242 0.9157 0.9104 0.9861 0.9817 0.9835 0.9507 0.0066
UDCNet TGRS’24 0.8821 0.8517 0.8214 0.9775 0.9689 0.9551 0.9311 0.0056 0.9239 0.9068 0.8932 0.9850 0.9803 0.9770 0.9497 0.0068
SOLNet TGRS’24 0.8609 0.8514 0.8394 0.9624 0.9576 0.9607 0.9171 0.0078 0.9012 0.8946 0.8925 0.9734 0.9688 0.9728 0.9284 0.0111
LiteSalNet TGRS’25 0.8878 0.8810 0.8798 0.9743 0.9633 0.9729 0.9276 0.0063 0.9124 0.9061 0.9056 0.9789 0.9699 0.9752 0.9372 0.0090
LGIPNet TGRS’25 0.8837 0.8702 0.8563 0.9761 0.9711 0.9701 0.9339 0.0065 0.9149 0.9027 0.8924 0.9806 0.9758 0.9755 0.9450 0.0082
DPU-Former IJCAI’25 0.8931 0.8758 0.8464 0.9816 0.9749 0.9665 0.9402 0.0056 0.9263 0.9142 0.9024 0.9868 0.9827 0.9838 0.9412 0.0062
Ours 0.9032 0.8876 0.8719 0.9801 0.9748 0.9457 0.9487 0.0049 0.9356 0.9243 0.9196 0.9861 0.9819 0.9650 0.9590 0.0056
@@ IV-A2 Implementation Details
Figure 6: The qualitative comparison between our method and six other approaches across various challenging scenarios.
IV-A3 Evaluation Metrics
@@
To comprehensively evaluate the performance of the proposed method, we adopt eight widely used evaluation metrics to assess the performance of the model. These metrics include the Mean Absolute Error ($MAE$), S-measure ($S_{\alpha}$), F-measure score ($F_{\beta}$), E-measure ($E_{\xi}$), adaptive E-measure ($E^{adp}_{\xi}$), mean E-measure ($E^{mean}_{\xi}$), maximum E-measure ($E^{max}_{\xi}$), adaptive F-measure ($F^{adp}_{\beta}$), mean F-measure ($F^{mean}_{\beta}$) and maximum F-measure ($F^{max}_{\beta}$). Specifically, $S_{\alpha}$ quantifies the structural similarity between the predicted saliency map and the ground truth. $F_{\beta}$ is the weighted harmonic mean of precision and recall, where $\beta^{2}$ is set to 0.3 to place greater emphasis on precision over recall. The $MAE$ calculates the average pixel-wise error between the prediction and the ground truth, offering a direct measure of prediction deviation. Furthermore, $E^{adp}_{\xi}$, $E^{mean}_{\xi}$, and $E^{max}_{\xi}$ evaluate the E-measure using adaptive thresholding, image-level mean, and maximum values, respectively. These metrics capture structural and distributional consistency from multiple perspectives. Similarly, $F^{adp}_{\beta}$, $F^{mean}_{\beta}$, and $F^{max}_{\beta}$ assess the F-measure under the same three thresholding strategies, emphasizing the trade-off between precision and recall.
IV-B Comparisons With State-of-the-Arts
@@
We compare our method with 15 state-of-the-art methods specifically designed for the ORSI-SOD task, including EMFINet[54], DAFNet[30], ERPNet[55], SBA-Net[56], HFANet[11], ESGNet[57], ACCoNet[40], BSCGNet[58], SFANet[10], RAGRNet[59], UDCNet[60], SOLNet[61], LiteSalNet[9], LGIPNet[62], and DPU-Former[25]. To ensure a fair evaluation, all experiments are conducted under the same conditions on the ORSSD and EORSSD datasets.
TABLE II: The effectiveness of each component in GCRPNet. In the tables, $\uparrow$ and $\downarrow$ indicate that higher and lower values are preferred, respectively.
Methods EORSSD ORSSD
w/o DS-HGAM 0.8907 0.9748 0.9389 0.0061 0.9135 0.9758 0.9416 0.0096
w/o MCAEM 0.8924 0.9752 0.9411 0.0055 0.9214 0.9835 0.9562 0.0059
w/o LESS2D 0.8870 0.9725 0.9394 0.0055 0.9170 0.9774 0.9450 0.0081
Ours 0.9032 0.9801 0.9487 0.0049 0.9356 0.9861 0.9590 0.0056
IV-B1 Quantitative Comparisons
@@
The quantitative comparison results between our proposed GCPRNet and other methods are summarized in Tab I. It is evident that our model achieves SOTA performances. On the EORSSD dataset, the quantitative comparison results of our method across eight evaluation metrics are presented at the bottom of the table. It can be observed that our approach achieves the best performance on several key indicators, including $F_{\beta}^{max}$, $F_{\beta}^{mean}$, $S_{\alpha}$, and $MAE$, while also attaining the second-best result on metrics $F^{adp}_{\beta}$, $E_{\xi}^{max}$, and $E_{\xi}^{mean}$ metric. Specifically, compared with the second-best method, our approach achieves performance gains of 0.82% on $F_{\beta}^{max}$, 0.66% on $F_{\beta}^{mean}$, and 0.85% on $S_{\alpha}$, further highlighting its superiority in salient object detection tasks. On the ORSSD dataset, our method also achieves superior performance. Specifically, it attains the best results on five key metrics: $F_{\beta}^{max}$, $F_{\beta}^{mean}$, $F_{\beta}^{adp}$, $S_{\alpha}$, and $MAE$, while also attaining the second-best results on metrics $E_{\xi}^{max}$ and $E_{\xi}^{mean}$. Compared to the second-best method, our approach achieves a 0.93% improvement on metric $F_{\beta}^{max}$, a 0.86% improvement on metric $F_{\beta}^{mean}$, and a 0.83% improvement on metric $S_{\alpha}$, while also reducing the $MAE$ by 0.06%, further validating the effectiveness and robustness of our approach.
@@
IV-B2 Quantitative Comparisons
@@
Figure 6 presents a qualitative comparison between our model and six of the latest state-of-the-art methods from recent years on the ORSSD and EORSSD datasets. We selected multiple representative ORSI scenarios, including large objects, small objects, elongated structures, and low-contrast objects. Compared with other methods shown in the figure, our model demonstrates superior performance in saliency detection under these complex conditions. For example, in the low-contrast scene in the seventh row, only our model is able to accurately segment the complete contour of the circular building, while other methods suffer from varying degrees of over-segmentation, indicating their limited adaptability to such conditions. This highlights our model’s strong capability in capturing subtle local variations and performing precise boundary localization. Similarly, in the fourth row, which involves small-object detection, our model correctly segments all three yellow objects with clear and non-overlapping boundaries. In contrast, other methods yield blurry and incomplete results, especially for the middle object. These results demonstrate our model’s excellent segmentation ability across different object scales, reflecting its powerful multi-scale feature representation. Moreover, in the l-shaped building scene in the last row, only our method and DPU-Former are able to fully recognize and segment the complete l-shaped structure, while all other methods fail to segment the complete building shape and exhibit obvious segmentation errors. In summary, the qualitative results in Figure 6 comprehensively demonstrate the robustness and accuracy of our proposed model in salient object detection across various complex ORSI scenes. Our approach not only ensures effective global context modeling but also excels at extracting critical local features, enabling the model to learn saliency information more effectively.
IV-C Ablation Studies
In this section, we conduct comprehensive ablation studies on the EORSSD and ORSSD datasets to demonstrate the effectiveness of each component in our model. The specific results are shown in Table II.
In the table, w/o DS-HGAM indicates that the difference-similarity guided hierarchical graph attention module is removed, meaning that encoder features are directly passed to the decoder via skip connections. w/o MCAEM denotes the removal of the multi-granularity collaborative attention enhancement module. w/o LESS2D indicates that the proposed locally enhanced 2D selective scan is not used in the decoder stage, and the conventional VMamba scanning strategy is adopted instead.
Our ablation study shows that the complete model achieves outstanding performance on both datasets. Specifically, it obtains scores of 0.9032, 0.9810, 0.9487, and 0.0049 on the four key metrics $F_{\beta}^{max}$, $E_{\xi}^{max}$, $S_{\alpha}$, and $MAE$ for the EORSSD dataset, and 0.9356, 0.9861, 0.9590, and 0.0056 for the ORSSD dataset. In contrast, the other variants exhibit varying degrees of performance degradation across these metrics. These results clearly demonstrate the effectiveness of our proposed modules. The experimental results of w/o DS-HGAM indicate that with the introduction of the DS-HGAM module, all metrics showed improvements. On the EORSSD dataset, metrics $F^{max}_{\beta}$, $E^{max}_{\xi}$, and $S_{\alpha}$ increased by 1.25%, 0.53%, and 0.98%, respectively, while metric $MAE$ decreased by 0.12%. Similarly, on the ORSSD dataset, metrics $F^{max}_{\beta}$, $E^{max}_{\xi}$, and $S_{\alpha}$ improved by 2.21%, 1.03%, and 1.74%, respectively, whereas metric $MAE$ decreased by 0.4%. These results indicate that introducing the DS-HGAM module into the skip connections helps the model to capture cross-spatial and cross-scale feature relationships, thereby enabling more effective feature integration and improving the quality of the saliency map. In the w/o MCAEM setting, the results demonstrate that with the introduction of the MCAEM module, metric $F^{max}_{\beta}$ on the EORSSD dataset increased from 0.8924 to 0.9032, achieving a performance improvement of 1.08%; metric $E^{max}_{\xi}$ improved from 0.9752 to 0.9801, with a gain of 0.49%, this demonstrates that multi-scale convolutions and attention mechanisms can effectively extract multi-scale local features, thereby enhancing the model’s ability to represent salient regions and significantly improving the quality of mask prediction. The experimental results of w/o LESS2D demonstrate that by using the LESS2D module, metric $F^{max}_{\beta}$ on the EORSSD dataset increased by 1.62%, metric $E^{max}_{\xi}$ increased by 0.76%, and metric $S_{\alpha}$ increased by 0.93%. On the ORSSD dataset, metric $F^{max}_{\beta}$ increased by 1.86%, metric $E^{max}_{\xi}$ increased by 0.87%, metric $S_{\alpha}$ increased by 1.4%, while metric $MAE$ decreased by 0.25%. The block-wise scanning strategy adopted by LESS2D enhances the dependencies among local pixels, facilitates the learning of fine-grained local feature representations, and fully leverages the sequence modeling capability of Mamba. This enables the model to more effectively integrate global context with local details, resulting in more precise saliency maps.
V Conclusion
In this paper, we propose a Mamba-based model for ORSI-SOD, named GCRPNet. Specifically, we design a DS-HGAM to perform interactive fusion of multi-scale features in the skip connections. By leveraging graph neural networks, DS-HGAM enables modeling of cross-spatial relationships to integrate features from different regions, thereby generating higher-quality saliency maps and enhancing the model’s robustness and generalization capability. Meanwhile, to address the limitations of existing Mamba models in local information modeling, we further propose a novel LEVSS Block and incorporate it into the decoding stage of the network. This module integrates two submodules: MCAEM and LESS2D. The MCAEM module expands the receptive field by employing depthwise separable convolutions with multiple kernel sizes and parallel attention mechanisms, enabling the model to better adapt to objects of varying scales and appearances. In addition, to further improve local modeling capability, the LESS2D module performs block-wise scanning on multi-scale feature maps, which strengthens local feature representations and effectively integrates global context modeling with local detail learning. Extensive experimental results demonstrate that the proposed model achieves state-of-the-art performance on ORSI-SOD tasks, fully validating its effectiveness and superiority.

### v4

@@ 3.5 Loss Function
4 Experiments
4.1 Experimental Setup
Datasets. To evaluate the effectiveness of the method, we trained and tested it on two public benchmark datasets for ORSI-SOD: ORSSD[23] and EORSSD[52].
@@ 4.1 Experimental Setup
Table 1: Quantitative comparison results in terms of $F_{\beta}^{max}$, $E_{\xi}^{max}$, $S_{\alpha}$ and MAEMAE score are presented on two benchmark datasets. In the tables, $\uparrow$ and $\downarrow$ indicate that higher and lower values are preferred, respectively. The best result in each column is highlighted in bold, and the second-best result is underlined.
Method Year EORSSD ORSSD
EMFINet 2021 .8721 .9712 .9292 .0084 .9002 .9737 .9366 .0109
ERPNet 2022 .8744 .9665 .9254 .0082 .9036 .9738 .9352 .0114
ESGNet 2023 .8907 .9665 .9374 .0051 .9220 .9858 .9484 .0065
BSCGNet 2023 .8809 .9693 .9296 .0079 .9101 .9784 .9411 .0093
ACCoNet 2023 .8822 .9759 .9303 .0067 .9149 .9819 .9428 .0087
SFANet 2024 .8834 .9770 .9351 .0058 .9192 .9830 .9453 .0077
RAGRNet 2024 .8853 .9785 .9364 .0057 .9242 .9861 .9507 .0066
UDCNet 2024 .8821 .9775 .9311 .0056 .9239 .9850 .9497 .0068
BCARNet 2025 .8871 .9762 .9361 .0054 .9196 .9833 .9465 .0071
MRBINet 2025 .8854 .9766 .9355 .0056 .9199 .9851 .9474 .0069
THMNet 2025 .8981 .9827 .9432 .0054 .9251 .9843 .9503 .0071
LGIPNet 2025 .8837 .9761 .9339 .0065 .9149 .9806 .9450 .0082
DPU-Former 2025 .8931 .9816 .9402 .0056 .9263 .9868 .9412 .0062
ASCNet (ours) .9009 .9838 .9449 .0048 .9383 .9924 .9582 .0058
@@
Implementation Details. In this study, we implemented the proposed method using the PyTorch framework and conducted all experiments on an RTX 4090. We set the learning rate to 1e-4, the batch size to 4, and employed the AdamW optimization strategy for training. To ensure consistency in the input data, each image was resized to 384×384 pixels before being fed into the model. Additionally, to effectively mitigate overfitting during training, we introduced a series of data augmentation techniques, including random scaling and cropping, random image transformations, and normalization. For the encoder, we adopted VMamba[30] initialized with publicly available pre-trained weights and fine-tuned it end-to-end on the target datasets. Finally, we trained the model for 100 epochs on the ORSSD and EORSSD datasets.
@@
Figure 5: The qualitative comparison between our method and other seven SOTA approaches across various challenging scenarios.
Figure 6: Quantitative comparison on the EORSSD and ORSSD datasets. (a) $F_{\beta}^{\max}$, (b) $E_{\xi}^{\max}$, (c) $S_{\alpha}$, and (d) MAE.
@@
Evaluation Metrics. To comprehensively evaluate the performance of the proposed method, we adopt four widely used evaluation metrics to assess the performance of the model. These metrics include the Mean Absolute Error ($MAE$), S-measure ($S_{\alpha}$) and maximum F-measure ($F^{max}_{\beta}$). Specifically, $S_{\alpha}$ quantifies the structural similarity between the predicted saliency map and the ground truth. $F_{\beta}$ is the weighted harmonic mean of precision and recall, where $\beta^{2}$ is set to 0.3 to place greater emphasis on precision over recall. The $MAE$ calculates the average pixel-wise error between the prediction and the ground truth, offering a direct measure of prediction deviation. Furthermore, $E^{max}_{\xi}$ evaluate the E-measure using maximum values. Similarly, $F^{max}_{\beta}$ assess the F-measure using maximum values, emphasizing the trade-off between precision and recall.
4.2 Performance Comparisons
@@
We compare our method with 13 state-of-the-art methods specifically designed for the ORSI-SOD task, including EMFINet[56], ERPNet[57], ESGNet[12], BSCGNet[9], ACCoNet[26], SFANet[34], RAGRNet[55], UDCNet[40], BCARNet[16], MRBINet[22], THMNet[49], LGIPNet[37], and DPU-Former[39]. To ensure a fair evaluation, we adopt the saliency maps generated using the official implementations or pretrained models provided by the original authors, and all methods are evaluated using the same evaluation toolkit and protocols on the ORSSD and EORSSD datasets
@@
Quantitative Comparisons. Figure 6 presents a visual quantitative comparison on the EORSSD and ORSSD datasets in terms of (a) $F_{\beta}^{\max}$, (b) $E_{\xi}^{\max}$, (c) $S_{\alpha}$, and (d) MAE. As shown in the figure, our method consistently achieves superior performance across all evaluation metrics on both datasets. The detailed quantitative results of our proposed ASCNet and other state-of-the-art methods are summarized in Tab. 1. It is evident that our model achieves SOTA performances. On the EORSSD dataset, the quantitative comparison results of our method across four evaluation metrics are presented at the bottom of the table. It can be observed that our approach achieves the best performance on all key indicators, including $F_{\beta}^{max}$, $E_{\xi}^{max}$, $S_{\alpha}$, and $MAE$. Specifically, compared with the second-best method, our approach achieves performance gains of 0.28% on $F_{\beta}^{max}$, 0.11% on $E_{\xi}^{max}$, and 0.17% on $S_{\alpha}$, further highlighting its superiority in salient object detection tasks. On the ORSSD dataset, our method also achieves superior performance. Specifically, it also achieved the best results across all metrics, compared to the second-best method, our approach achieves a 1.2% improvement on metric $F_{\beta}^{max}$, a 0.56% improvement on metric $E_{\xi}^{max}$, and a 0.79% improvement on metric $S_{\alpha}$, further validating the effectiveness and robustness of our approach.
@@
Table 2: Comparison of different module ablation settings on ORSSD and EORSSD datasets. The best result in each column is highlighted in bold.
NO. Method ORSSD EORSSD
MLCM GPM DAGS $S_{\alpha}\!\uparrow$ $F_{\beta}^{max}\!\uparrow$ $E_{\xi}^{max}\!\uparrow$ $MAE\!\downarrow$ $S_{\alpha}\!\uparrow$ $F_{\beta}^{max}\!\uparrow$ $E_{\xi}^{max}\!\uparrow$ $MAE\!\downarrow$
(a) .9049 .879 .9663 .0178 .9078 .8417 .9566 .0100
(b) .9284 .8981 .9761 .0121 .9249 .8681 .9698 .0070
(c) .9397 .9119 .9794 .0090 .9354 .8853 .9710 .0052
(d) .9326 .9058 .9778 .0107 .9310 .8797 .9758 .0060
(e) .9530 .9273 .9865 .0062 .9421 .8972 .9831 .0056
(f) .9474 .9275 .9860 .0072 .9329 .8862 .9782 .0052
(g) .9434 .9220 .9849 .0078 .9404 .8956 .9822 .0050
Ours .9582 .9383 .9924 .0058 .9449 .9009 .9838 .0048
Figure 7: Visual comparison of qualitative results produced by models with different ablation settings.
@@
Qualitative Comparisons. Figure 5 presents a qualitative comparison between our model and six of the latest state-of-the-art methods from recent years on the ORSSD and EORSSD datasets. We selected multiple representative ORSIs scenarios, including large objects, small objects, elongated structures, and low-contrast objects. Compared with other methods shown in the figure, our model demonstrates superior performance in saliency detection under these complex conditions. For example, in the low-contrast scene in the fourth row, only our model is able to accurately segment the complete contour of the ship, whereas other methods fail to correctly identify the low-contrast ship and exhibit varying degrees of segmentation errors, indicating their limited adaptability to such challenging conditions. This highlights our model’s strong capability in capturing subtle local variations and performing precise boundary localization. Similarly, in the fifth row, which involves elongated structures detection, only our model is able to detect the complete contour of the target, whereas the other methods are almost unable to effectively recognize it. This result clearly demonstrates that the proposed model, while fully leveraging the long-range dependency modeling capability of the state space architecture, also exhibits excellent fine-grained boundary localization and structural perception abilities. Moreover, in the small objects scenario shown in the last row, only our method able to accurately detect all salient objects, whereas the other methods generally suffer from missed detections. This phenomenon further indicates that our model possesses stronger target discrimination capability and scale robustness in complex multi-scale scenes. In summary, the qualitative results in Figure 5 comprehensively demonstrate the robustness and accuracy of our proposed model in salient object detection across various complex ORSIs scenes. Our approach not only ensures effective global context modeling but also excels at extracting critical local features, enabling the model to learn saliency information more effectively.
4.3 Ablation Studies
Effectiveness of Key Components. We evaluate the effectiveness of the proposed MLCM, GPM, and DAGS through a series of ablation experiments on the ORSSD and EORSSD datasets, as summarized in Table 2. Starting from a baseline model without any proposed modules, introducing each component individually leads to consistent performance improvements. Specifically, MLCM enhances global structural perception via multi-level context aggregation, GPM yields more pronounced gains by reinforcing global contextual priors and suppressing background interference, while DAGS further improves boundary completeness and structural coherence through resolution-aware sequence modeling. Combining any two modules consistently outperforms their single-module counterparts, indicating strong complementarity among them. The best performance is achieved when all three modules are jointly employed, yielding the highest scores across all evaluation metrics on both datasets. The qualitative results in Figure 7 further support these findings: the baseline model tends to produce incomplete predictions with fragmented boundaries and background confusion, while the full model generates more coherent object regions with sharper boundaries and fewer false positives, especially in challenging scenarios with complex backgrounds and elongated structures. These results demonstrate that MLCM, GPM, and DAGS are highly complementary and collectively contribute to robust and accurate salient object detection.
DAGS with other scanning strategies. Table 3 compares the proposed DAGS with several representative scanning strategies on the EORSSD dataset. As shown, DAGS consistently outperforms Cross Scan[30], Diagonal Scan[45], Hilbert Scan[17], and Local Scan[21] across all metrics, achieving the highest $F_{\beta}^{max}$, $E_{\xi}^{max}$, and $S_{\alpha}$, while yielding the lowest MAE. These results demonstrate that the proposed DAGS is more effective in capturing spatial dependencies and structural information, thereby validating the superiority of our scanning strategy.
Table 3: Ablation for the impact of different scanning.
ID Scanning Method EORSSD
$F_{\beta}^{max}\uparrow$ $E_{\xi}^{max}\uparrow$ $S_{\alpha}\uparrow$ $MAE\downarrow$
1 Cross Scan [30] .9421 .8972 .9831 .0056
2 Diagonal Scan [45] .9391 .8917 .9782 .0064
3 Hilbert Scan [17] .9384 .8906 .9755 .0057
4 Local Scan [21] .9407 .8947 .9812 .0059
5 DAGS (Ours) .9449 .9009 .9838 .0048
5 Conclusion
In this paper, we propose ASCNet, a novel network for ORSI-SOD, designed to address the challenges of complex backgrounds, large-scale variations, and intricate spatial structures in optical remote sensing images. Specifically, we introduce a MLCM to enable interactive fusion of multi-scale features across different representation levels. By leveraging topology-aware attention to model cross-spatial and cross-regional relationships, MLCM effectively aggregates contextual information from multiple semantic levels and spatial regions, resulting in more discriminative saliency representations with improved robustness and generalization. To enhance local region modeling in sequence-based visual representations, we further integrate an APVSS block into the decoder, which consists of GPM and DAGS. The GPM employs learnable global semantic tokens to semantically condition local patch features, strengthening local structural perception while maintaining global semantic consistency. Meanwhile, DAGS adopts a resolution-aware, block-wise scanning strategy on multi-scale feature maps, adjusting scan granularity by resolution and dynamically reweighting directional scans to balance local detail modeling and global contextual awareness during decoding. Extensive experiments on multiple ORSI-SOD benchmarks demonstrate that ASCNet consistently achieves state-of-the-art performance, validating the effectiveness of the proposed modules and overall architecture.